I need your help with this: Simplify the power of i,1,-i or -1

i71

Answer by stanbon(75887)

I need your help with this: Simplify the power of i,1,-i or -1
i^(71) = ?
---
Note: 71 = 4*17+3
= i^(4*17+3)
---
= [i^4]^17 * i^3
---
= 1^17 *(-i)
---
= 1 * -i
--------------
= -i
-----
Cheers,
Stan H.
========